At ABC University, a group of students was selected and asked how much of their weekly allowance they spent in buying mobile phone load. The following is the list of amounts spent: Php 120, 110, 100, 200, 10, 90, 100, 100. find the standard deviation.



1
Expert's answer
2021-09-27T16:17:12-0400

Formula for the standart deviation(unbiased): "s^2=\Sigma (x(i)-u)^2/(n-1)", where x(i)-element with i number, u-mean of the sample, n-sample size.

Mean of the sample: "u=(120+...+100)\/8=103.75"

Sum:

"\Sigma (x(i)-u)^2=(120-103.75)^2+...+(100-103.75)^2= 18413.6875"

"s^2=18413.6875\/7=2630.5(approximately)"

standart deviation is equal to s, and approximately equal to 51.3